# Function-Space MCMC for Bayesian Wide Neural Networks

### v1

@@ 5 Conclusion
In this paper, we investigated the effectiveness of the pCN and pCNL samplers in sampling wide Bayesian Neural Networks (BNNs), and we provided insights into the scalability and efficiency of these methods; particularly when compared to the underdamped LMC sampler. Our theoretical findings, grounded in the understanding of the reparameterized posterior distribution, demonstrated that the acceptance probability of the pCN proposal converges to $1$ as the network width goes to infinity, thus ensuring efficient sampling. Empirical results further corroborated this theoretical prediction. When applied to fully connected networks with varying widths, pCN consistently demonstrated higher per-step ESS compared to LMC, especially as dimensionality increased. This finding is crucial as it underscores the advantage of the pCN approach in wide Bayesian neural networks, where standard MCMC techniques often struggle due to high autocorrelation among samples. Our results indicate that the pCN sampler, under the proposed framework, offers a reliable and scalable solution for Bayesian inference in wide neural networks, ultimately leading to better diagnostic performance and improved computational efficiency. The dimensional robustness of the pCNL method, along with the incorporation of gradient information within the MH proposal, makes it effective and stable for posterior sampling from BNNs. However, the method is significantly more expensive computationally compared to pCN and therefore, given the absence of significant increases in performance as the network width grows larger, the pCN version is preferred in very large width regimes.
Limitations
Due to constraints in our computational resources, we were not able to run experiments on BNNs with wider layers. However, based on our current analyses, as presented in the acceptance probability analysis of Figure 2 and in the ESS analysis of Figure 3, we anticipate a widening performance gap between the LMC and pCN samplers with increased layer width. Theoretically, we showed that the acceptance rates for pCN and pCNL converge to $1$, supporting the scalability and robustness of the algorithms. In contrast, the LMC sampler’s acceptance rate is known to decrease further as layer width increases. This anticipated behavior suggests an increase in autocorrelation among samples from the LMC, and a decrease in autocorrelation for the pCN samples. These projections underscore the potential limitations of LMC samplers in handling larger BNNs, highlighting the need for more robust sampling methods in extensive computational settings. Finally, it would be interesting to extend these empirical analyses to other neural network architectures and to deep neural networks.
Appendix A Proof of Theorem 2.1
@@
We have already shown that $q(\phi^{*}|\phi)=\mathcal{N}(\sqrt{1-\beta^{2}}\phi,\beta^{2}I_{D})$. Regarding the reparamatrized weight posterior of the network, we observe that [15]
@@
where $p(\phi^{(L+1)}|\phi^{(\leq L)},\mathcal{D})\sim\mathcal{N}(0,I_{d^{(L)}})$ is assured by the reparametrisation. It is then crucial to recognize the empirical NNGP kernel $\hat{K}_{\sigma^{2}}=\sigma^{2}I_{n}+\Psi\Psi^{T}$ [25] and observe that $det(\Sigma)\propto det(\hat{K}_{\sigma^{2}})$.
@@
Let’s start by writing explicitely the expressions of the acceptance probability for the pCNL algorithm:
@@
By definition 2, $p(W|\mathcal{D})=p(\phi^{(\leq L)}|\mathcal{D})$ and thus we can exploit the known expressions from Appendix A:
@@
That corresponds to the same exact expression found in Appendix A and hence convergence to 1 as the layers width increases is granted by the same arguments.

### v4

@@ 5 Discussion
We investigated the pCN and pCNL algorithms in sampling wide BNNs, providing insights into their scalability and efficiency, in comparison to the underdamped LMC sampler. Our theoretical findings, grounded in the understanding of the reparameterized posterior distribution, demonstrated that the acceptance probability of the pCN and pCNL proposals converges to $1$ as the network width goes to infinity, thus ensuring efficient sampling. Empirical results further corroborated this theoretical prediction, showing that, when applied to fully connected networks with large widths, these function-space MCMC algorithms consistently achieved higher per-step ESS than LMC, especially as dimensionality increased. This finding is significant as it underscores the advantage of these approaches in wide BNNs, where standard MCMCs often struggle due to non-robustness and high autocorrelation among samples.
More precisely, our results indicate that the pCN sampler, within the proposed framework, offers a reliable and scalable solution for Bayesian inference in the context of wide neural networks, ultimately leading to better diagnostic performance, as well as an improved computational efficiency. However, for neural networks with smaller widths, the pCN method appears too naive, and it is significantly outperformed by its Langevin counterpart.
Both the pCN and the pCNL show promise as complementary alternatives to standard MCMC algorithms for wide BNNs. While the pCN algorithm is scalable and highly effective in very large networks, its performance decreases sharply as the number of neurons decreases. The gradient information within the MH proposal shows potential as a solution to this limitation, yielding more stable and meaningful results even in lower-width regimes. However, pCNL is more computationally expensive and does not offer clear advantages in large-scale settings, making it less preferable in such cases.
5.1 Future work
While our study provides valuable insights into the effectiveness of the pCN and pCNL algorithms in sampling from wide BNNs, it is important to acknowledge several limitations that may affect the generalizability of our findings.
We limited our experiments to one-layer fully connected BNNs in order to isolate and highlight the impact of width changes on the performance of the samplers. More extensive experiments with deeper networks and different architectures are necessary to fully explore the applicability of these methods in real-world case studies. Testing the algorithms on a wider variety of datasets will help determine their robustness across diverse tasks.
In terms of future works, an interesting direction involves the analysis of our methods in conjunction with existing studies on the mixing speed of MCMC algorithms. Specifically, a key challenge in the field is to establish that the number of steps required for the marginal distribution of MCMC procedures to converge to the target distribution does not grow exponentially with the number of parameters. For the pCN algorithm, this question has been addressed positively in Theorem 2.12 of Hairer et al., (2014), where conditions under which the number of steps remains manageable are provided. Since, loosely speaking, the main assumption requires the acceptance rate to have a lower bound that is strictly positive, an application of the result in our framework seems possible. Indeed, in the context of our work, we have this assumption for granted by showing that the acceptance rate converges to 1 as the network width increases. Thus, it would be interesting to combine our results with these findings to explore further implications for chain mixing properties, potentially offering new insights into the efficiency of pCN and pCNL in high-dimensional settings. See also Andrieu et al., (2024).
Acknowledgments
S.F. acknowledges the Italian Ministry of Education, University and Research (MIUR), “Dipartimenti di Eccellenza” grant 2023-2027.
Appendix A Proofs and derivations
A.1 Proof of Theorem 2.1
@@ Appendix A Proofs and derivations
We have already shown that $q(\phi^{*}|\phi)=\mathcal{N}(\sqrt{1-\beta^{2}}\phi,\beta^{2}I_{D})$. Regarding the reparamatrized weight posterior of the network, we observe that Hron et al., (2022)
@@
where $p(\phi^{(L+1)}|\phi^{(\leq L)},\mathcal{D})\sim\mathcal{N}(0,I_{d^{(L)}})$ is assured by the reparametrisation. It is then crucial to recognize the empirical NNGP kernel $\hat{K}_{\sigma^{2}}=\sigma^{2}I_{n}+\Psi\Psi^{T}$ (Rasmussen and Williams,, 2006) and observe that $det(\Sigma)\propto det(\hat{K}_{\sigma^{2}})$.
@@
Let’s start by writing expliciteply the expressions of the acceptance probability for the pCNL algorithm:
@@
By definition 1, $p(W|\mathcal{D})=p(\phi^{(\leq L)}|\mathcal{D})$ and thus we can exploit the known expressions from Section A.1:
@@
That corresponds to the same exact expression found in Section A.1 and hence convergence to 1 as the layers width increases is granted by the same arguments.
Appendix B Additional experiments
B.1 Gelman-Rubin Statistic Analysis
The Gelman-Rubin statistic or potential scale reduction factor, $\hat{R}$, (Gelman and Rubin,, 1992; Brooks and Gelman,, 1998) complements the ESS analsys by assessing how the different samples affect the rate of convergence of the collected Markov chain. More precisely, the metric is evaluated by running multiple independent chains per each sample and then exploiting comparisons of the variance between multiple chains to the variances within the single chains to give a measure of the degree to which variance (of the means) between chains exceeds what one would expect if the chains were identically distributed. Mathematically, the Gelman-Rubin statistic is defined as follows: Suppose we have run $M$ chains in parallel with different starting values and that we have collected $N$ samples for each chain. Denote by $x_{i}^{(m)}$ the $i$-th sample of the $m$-th chain. The within-chain variance and the between-chain variance are defined as:
$\hat{R}=\frac{\frac{N-1}{N}W+\frac{1}{N}B}{W}.$
Now, since at convergence all the parallel chains are indistinguishable, the Gelman-Rubin statistic should be close to 1. In practice, a value of $\hat{R}$ smaller than $1.2$ is considered as a good indicator of approximate convergence (Gelman and Rubin,, 1992).
Figure 3: Evolution of the Gelman-Rubin statistic, evaluated using three independent chains, of the LMC, pCN and pCNL as a function of the number of steps for the stepsize $\beta=0.1$. The solid lines represent the average Gelman-Rubin statistic, whereas the shaded areas indicate their standard deviation. Again, the CIFAR-10 dataset, with sample size fixed at $n=256$, is used. Since the metric should be close to 1 for all chains to be considered converged, we trace a horizontal line at $\hat{R}=1.2$, indicating the standard empirical threshold for determining convergence. For all samples the chosen burn-in of 20,000 steps appears to be sufficient to ensure that the chains have reached the target distribution. The metric for pCN sampler improves as the width increases, whereas the LMC method shows complementary results. Finally, the pCNL approach exhibits consistent and good results among all dimensions.
We examine the $\hat{R}$ values for the pCN, LMC, and pCNL samplers, considering their evolution as a function of the number of steps [5000, 10000, 15000, 20000]. The stepsize is fixed at $\beta=0.1$. For each sampler, we run three independent chains for each of the BNN widths among $\{512,1024,2096,4192\}$, we then use the collected samples to calculate the metric values. More specifically, our focus is on investigating how the $\hat{R}$ values for different widths of the BNN change as the number of steps increases for the different samplers. This analysis allows us to evaluate and compare the convergence efficiency of the different samplers across these varying network sizes. The results are presented in Figure 3, where we also plot the reference line of $\hat{R}=1.2$ to indicate the standard empirical threshold for determining convergence.
B.2 Trace plots of iterates PCs
To further assess the convergence behavior of the MCMC samplers, we analyze the trace plots of the principal components (PCs) of the collected iterates. These plots visualize the evolution of the dominant modes of variation in the sampled posterior over time, providing insight into whether the Markov chain has reached stationarity. Specifically, well-mixed chains should exhibit stable oscillations around a mean value, without strong trends or prolonged correlations. Below, we present the trace plots of the first two principal components for all the considered sampling methods and stepsizes.
Figure 4: Trace plots of the first two principal components for the LMC sampler with different stepsizes. From top to bottom: $\beta=0.01$, $\beta=0.1$, and $\beta=0.2$. A well-mixed chain should exhibit stable oscillations around a mean value, without strong trends or correlations.
We can notice in Figure 4 how the LMC sampler mixing properties deteriorate with the layers width, reflecting its well-known sensitivity to high-dimensional settings.
Figure 5: Trace plots of the first two principal components for the pCN sampler with different stepsizes. From top to bottom: $\beta=0.01$, $\beta=0.1$, and $\beta=0.2$. The stability of the trace plots improves for larger network widths, demonstrating the suitability of pCN in wide BNNs.
The pCN sampler, Figure 5, shows the opposite behavior: as the network width increases, the trace plots become more stable, supporting the theoretical guarantee that its acceptance rate converges to 1 in the infinite-width regime.
Finally, in Figure 6, appears that the pCNL samples maintains good mixing properties also in the smaller width regimes.
Figure 6: Trace plots of the first two principal components for the pCNL sampler with different stepsizes. From top to bottom: $\beta=0.01$, $\beta=0.1$, and $\beta=0.2$. The performance of pCNL is analyzed to understand its convergence properties in comparison to LMC and pCN.
Appendix C Further mathematical foundations and justifications
C.1 Fréchet derivative
The Fréchet derivative is a generalization of the derivative to function spaces. Given a Banach space $X$ and a function $F:X\to Y$ mapping between Banach spaces, we say that F is Fréchet differentiable at a point $\theta\in X$ if there is a bounded linear operator $DF(\theta):X\to Y$ satisfying:
$\lim_{\|h\|\to 0}\frac{\|F(\theta+h)-F(\theta)-DF(\theta)[h]\|_{Y}}{\|h\|_{X}}=0.$ (10)
We call $DF(\theta)$ the Fréchet derivative of F at $\theta$.
This definition ensures that $DF(\theta)$ provides the best linear approximation of $F$ around $\theta$, analogous to the Jacobian in finite-dimensional spaces.
The Fréchet derivative of the log-likelihood plays a crucial role in the definition of the pCNL sampler, as it directly influences the proposal dynamics. Here we add details on its computation. Given the likelihood function:
$p(y|\theta,x)\sim\mathcal{N}(f^{(L+1)}(x),\sigma^{2}I),$ (11)
the corresponding log-likelihood function is:
$\ell(\theta)=-\frac{n}{2}\log(2\pi\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\|y_{i}-f^{(L+1)}(x_{i})\|^{2}.$ (12)
To compute the Fréchet derivative, we differentiate $\ell(\theta)$ with respect to $\theta$. Since $f^{(L+1)}(x)$ depends on the network parameters $\theta$, we apply the chain rule:
$D\ell(\theta)=-\frac{1}{\sigma^{2}}\sum_{i=1}^{n}(y_{i}-f^{(L+1)}(x_{i}))\nabla_{\theta}f^{(L+1)}(x_{i}),$ (13)
where $\nabla_{\theta}f^{(L+1)}(x)$ is the Jacobian of the network output with respect to the parameters $\theta$, computed recursively using backpropagation.